Find the vertex of y = |2x+4|-3
Answer by stanbon(75887)

Find the vertex of y = |2x+4|-3
---
The turn-around point is where 2x+4 = 0
Solve 2x+4=0
2x = -4
x = -2
---
The corresponding y value is:
f(-2) = 0-3
----
Vertex at (-2,-3)
======================
Cheers,
Stan H.